A person attempting to maintain social distancing is walking at 1.20 m/s and is being overtaken by a person walking at 1.58 m/s. If the faster person starts 13.3 m behind the slower person, how long does it take for the faster person to come within 2.00 m of the slower person?
Answer #1

Speed of the person attempting to maintain social distancing = V1 = 1.2 m/s

Speed of the faster person = V2 = 1.58 m/s

Initial distance of the faster person from the slower person = D1 = 13.3 m

Final distance of the faster person from the slower person = D2 = 2 m

Distance the faster person covers more than the slower person = D

D = D1 - D2

D = 13.3 - 2

D = 11.3 m

Time taken by the faster person to come within 2 m of the slower person = T

Distance traveled by the slower person in this time = L1

L1 = V1T

Distance traveled by the faster person in this time = L2

L2 = V2T

D = L2 - L1

D = V2T - V1T

D = (V2 - V1)T

11.3 = (1.58 - 1.2)T

T = 29.7 sec

Time taken by the faster person to come within 2 m of the slower person = 29.7 sec